1. A researcher uses a simple linear regression to measure the relationship between the monthly salary (Salary measured in dollars) of data scientists and the number of years since being awarded a Master degree (Master Degree). A random sample of 80 observations was collected for the analysis. A researcher used the econometric model which has the following specification Salary,-β0 + β, Master-Degree, + εί, where i = 1, , 80 The (incomplete) Excel output of equation (1) is presented below SUMMARY OUTPUT Regression Statistics R-Square Observations ANOVA XxX 80 df MS Regression Residual Total XxX 126408.2 XxX 20514650.04 Standard Coefficients error t-stat Intercept Master_Degree 6589.12 150.95 136.14 13.52

a) What are the assumptions on the error term (Ei) a researcher has made in the (3 marks) (2 marks) simple linear regression (1)? b) State the estimated equation.

c) Interpret the slope coefficient. Is the slope coefficient in line with your expectation? (3 marks) (4 marks) e) Perform a test to determine whether there is a significant positive relationship d) Determine the 90% confidence interval for the slope coefficient. between the monthly salary and the number of years since being awarded a Master degree (at the 5% level of significance). (4 marks) f Calculate and interpret the coefficient of determination. g) Predict the expected monthly salary of a data scientist who has been awarded with a master degree for 20 years. (3 marks)

Simple Linear regression
Answer #1

The regression model being estimated is

Salary Bo+B1Master Degree, +e,

where Salary is the Monthly salary (in Dollars)

Master_Degree is the number of years since being awarded a master degree.

a) The error term is epsilon_istackrel{iid}sim mathcal{N}(0,sigma^2)

The assumptions are

  • The error term is i.i.d, that is the errors are independent and identically distributed.
  • The error terms are normally distributed with mean 0 and constant variance

b) The estimated values of the coefficients are

Coefficients Intercept Master_Degree 6589.12 150.95

The estimated value of intercept is 30 = 6589.12

The estimated value of the slope is 31 = 150.95

Ans: The estimated equation is

Salary 6589.12 + 150.95 × Master-Degree

c) The estimated value of the slope coefficient is 150.95. The positive value indicates that number of years since being awarded a master degree and the monthly salary move in the same direction. That is the monthly salary would increase with the increase in number of years since being awarded a master degree. For each year increase in the number of years since being awarded a master degree, the monthly salary increases by $150.95.

It is reasonable that monthly salary would increase with the increase in number of years since being awarded a master degree, provided the person is gainfully employed (as a data scientist?) during these years in the field of interest. That is there is a corresponding increase in the number of years of experience working as a data scientist.

d) 90% confidence interval indicates that a significance level of a 1-90/100 = 0.10 .

The critical t value is obtained using P(T > ta/2) = a/2 = 0.1/2 = 0.05 .

The number of observations is n=80. The degrees of freedom for t is n-2=80-2=78.

Using the t tables we can get the critical value of t for degrees of freedom df=60 as to/2 = 1.671 and for df=120 as 1.658.

the value for df=78 will be something between these 2 values. We can either interpolate of use excel to get the exact value.

Using =T.INV.2T(0.1,78) we get a t value of 1.665

to/2 = 1.665

We know the standard error of the slope estimate from the output

se(31) 13.52

The 90% confidence interval is

egin{align*} &hat{eta}_1pm t_{alpha/2}s.e(hat{eta}_1) implies &150.95pm 1.665 imes 13.52 implies &[128.44,173.46] end{align*}

ans: 90% confidence interval for the slope is [128.44,173.46]

e) There would be a positive relationship between monthly salary and number of years, if the slope coefficient is positive (that is >0).

That is we want to test the following hypotheses

Ho: B1 -0 null hypothesis: There is no positive relationship between salary and number of years Ha : β1 > 0 ← alternative hypothesis: There is a positive relationship between salary and number of years a 0.05 <-level of significance to test the hypotheses

The hypothesized value of the slope is egin{align*} eta_{1H_0}=0 end{align*}

The test statistics is

1Ho =11.165 s.e Bi 13.52

this is a 1 tailed (right tailed) test (The alternative hypothesis has ">")

the p-value is P(T>11.165).

Using the excel function, =T.DIST.RT(11.165,78) we get p-value=3.83E-18

We will reject the null hypothesis if the p-value is less than the significance level.

Here the p-value of 0.000 is less than the significance level 0.05.

Hence we reject the null hypothesis.

We conclude that there is a significant positive relationship between monthly salary and number of years.

f) Using the ANOVA table

df MS Regression Residual Total 126408.2 20514650.04

the degrees of freedom for Residuals is df=n-2=80-2=78

The Mean square residuals is MSE=126408.2.

The sum of square residuals is

egin{align*}& ext{MSE}=rac{ ext{SSE}}{df} implies & ext{SSE}= ext{MSE} imes df implies & ext{SSE}=126408.2 imes 78=9859839.6 end{align*}

Sum of square Total is SST=20514650.04

The coefficient of determination is

egin{align*}R^2=1-rac{ ext{SSE}}{ ext{SST}}=rac{9859839.6}{20514650.04}=0.5194 end{align*}

The value of coefficient of determination is 0.5194. It indicates that 51.94% of variation is monthly salary is explained by the variation is the number of years since being awarded a master degree.

g) The expected value of salary for master_degree=20 is

widehat{ ext{Salary}}=6589.12+150.95 imes 20=9608.12

The expected monthly salary of a data scientist who has been awarded with a master degree for 20 years is $9,608.12